n each of the following snerios, categories the price elasticity of demand as either elastic, inelastic or unit elastic.(use mid point method in calculation)



a) The price of personal computers falls from $2750 to $2250 and the quantity demanded increases from 40,000 to 60,000



b) Pakistan posts increases the price of stamp for Rs.2 cents to Rs 4,but it's total revenue unchanged

1
Expert's answer
2021-12-07T10:17:19-0500

#Task 275621

Solution:

Mid point formula:

"\frac{Q2-Q1}{\frac{Q2+Q1}{2}}÷\frac{P2-P1}{\frac{P2+P1}{2}}"

"P1=\$2750\\P2=\$2250\\Q1=40000\\Q2=60000"

"\frac{60000-40000}{\frac{60000+40000}{2}}÷\frac{2250-2750}{\frac{2250+2750}{2}}\\\frac{20000}{50000}÷\frac{-500}{2500}=-2"

Taking absolute value

"Ep=2"


Indicating the price elasticity of demand is elastic


(b) An increase in price brings the same total revenue is unit elastic demand which is a type of demand which changes in the same proportion to its price; this means that the percentage change in demand is exactly equal to the percentage change in price.